Complete the following problems from Chapters 4 and 5 in Managerial Economics: A Problem Solving Approach. For each question, write 250-500 word explanation clearly showing how you solved the problem.

  1. Chapter 5: 5-5

Last year, a toy manufacturer introduced a new toy truck that was a huge success. The company invested $2.5 million for a plastic injection molding machine (which can be sold for $2.0 million) and $100,000 in plastic injection molds specifically for the toy (not valuable to anyone else).Labor and the cost of materials necessary to make each truck is about $3.This year, a competitor has developed a similar toy that has significantly reduced demand for the toy truck .Now, the original manufacturer is deciding whether they should continue production of the toy truck. If the estimated demand is 100, 000 trucks, what is the break-even price for the toy truck? Should you shut down?
Answer #1

Break even point is the point where AVC (Average Variable Cost) = MR (Marginal revenue).

Here, Total Variable Cost = Total Revenue.

That is, 100000*3 = 100000*P => Break even price, P = $3

Shut down point is the point where ATC(Average Total Cost) = MR (Marginal revenue).

Here, Total cost = Total revenue

Total cost = Total variable cost + Total fixed cost = 100000*3 + 100000 + 2500000 = $2900000

At break even price, Profit = TR - TC = 300000 - 2900000 = -$2600000 (loss)

Although the firm is suffering overall loss, it is still able to recover its variable cost at break even point. It should not shut down at P=3.